A loan of $1400 is due in 1 year at a simple interest rate of 12%. Partial payments of $400 in 2 months, $25 in 6 months, and $700 in 8 months are made

A. Construct a timeline clearly showing the timing of the debt and all repayments.

B. Determine the balance due at the end of the 12 months using the merchant rule

C. Determine the balance due at the end of the 8 months using the declining balance method
1
Expert's answer
2020-05-11T10:46:41-0400

A.2 month - 400, 6 month - 25, 8 month - 700,

rest - 373.5

B."400\times(1+0.12\times\frac{10}{12})+25\times(1+0.12\times\frac{6}{12})+700\times(1+0.12\times\frac{4}{12})+x=1400\times(1+0.12\times\frac{12}{12})"

440+26.5+728+x=1568

x=373.5

C.Balance due in two months:

"1400\times(1+0.12\times\frac{2}{12})-400=1028"

Interest due in four months:

"1028\times 0.12\frac{4}{12}=41.12"

41.12>25

Balance due in six months:

"1028\times(1+0.12\times\frac{6}{12})-(25+700)=1089.68-725=364.68"